- Use the Factor Theorem to show that x- cis a factor of Px) for the given values of c. P(x) = 2x0 - 17x2 - 10x + 9; c=
- Write the equation of a parabola that has a vertex at (-3, 3) and a directrix at x = 3
Answer #1

Solution Part I Griven that Plat) = 2x3 - Tablox to jente and not as factor of Pon, use factor Shorum loremaunder the arom) PHere a=6 (chk) = (33) eqn of parabola - 609 4 = 9234) - 4914 k) = (arch? 4x6 14+3)= -(0-3)? [241 443) - - (2-3)? This is ooh